a) Consider the following information for company XYZ:


Current dividend = 4n


Dividend 2years ago= 3.3n


Current equity beta=1.6


Equity risk premium=10%


Risk free rate=5%



Calculate the price of a share for this company. [5marks]

1
Expert's answer
2022-03-07T11:04:14-0500

"Expected return = risk-free rate+beta(market risk premium)"

"ER = 5 +1.6(10)"

"ER = 21percent"

"Dividend growth rate = \frac{4+3.3}{2}"

=3.65%

"share value = \frac{expected annual dividend}{expected rate of return-dividend growth rate}"

"share value =\frac{4n}{21-3.65}"


"share value = 23.05n"